Calculate Vr, Hr, and Sr for the following using appropriate generalized correlation:
Pure CO2 at 400k and 200 bar.
A mixture of 40% ethane and 60% ethylene.
1
Expert's answer
2020-02-27T09:02:38-0500

For pure CO2 400K 200 bar

Tpc = 304.2 K

ppc = 73.83 bar

Tpr = 400/304.2 = 1.31

ppr = 200/73.83 = 2.71

At these Tpr and ppr :

Z0 = 0.659

Z1 = 0.207

(Hr/(RTpc))0 = -2.03

(Hr/(RTpc))1 = -0.243

(Sr/(R))0 = -1.15

(Sr/(R))1 = -0.413

ω=0.224

Z = Z0 +ω *Z1 = 0.705

Vr = RT/P*(Z-1) = 8.314* 400/(200*10^5) * ( 0.705-1) = -49.1 cm3/mol

(Hr/(RTpc)) = (Hr/(RTpc))0 + ω *(Hr/(RTpc))1 = -2.08

Hr = -2.08 * 8.314 * 304.2 = -5261 J/mol

(Sr/(R)) = (Sr/(R))0 + ω *(SR/(R))1 = -1.24

Sr = -1.24 * 8.314 = -10.3 J/(mol*K)


all reference data was taken from

INTRODUCTION TO CHEMICAL ENGINEERING THERMODYNAMICS

J. M. Smith